OpenStudy (anonymous):

The midpoint of line segment AB is M. If the coordinates of A are (2,-6) and the coordinates of M are (5,-1), find the coordinates of B.

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

-2+2/2, 4+(-2)

OpenStudy (anonymous):

(0,3.5 or 4)

hartnn (hartnn):

(2+x)/2 = 5 find x from here

hartnn (hartnn):

(-6+y)/2 = -1 find y from here

OpenStudy (anonymous):

SO the coordinates of B are not 0,4?

hartnn (hartnn):

no

OpenStudy (anonymous):

Is x=5

hartnn (hartnn):

(2+x)/2 = 5 2+x = 10 x = ...

OpenStudy (anonymous):

x=5

hartnn (hartnn):

2+x = 10 x= 8

OpenStudy (anonymous):

You subtract 2 from both sides?

hartnn (hartnn):

yes

OpenStudy (anonymous):

Now do the same with y, right?

OpenStudy (anonymous):

@d3v13

OpenStudy (anonymous):

(8,4)

OpenStudy (anonymous):

y=4?

hartnn (hartnn):

(-6+y)/2 = -1 -6+y = -2 y = -2+6 =4